27 test scores in a math class have a standard deviation of 12.3. use a .04 significance level to test the claim that this class has a standard deviation less than 14.1

Step 1. H0:

ha:

Step 2:significance level a= .05

step 3: test (Please state what kind of test used. )

values for each required entry on calculator! (Please State)

Step 4: decision :

Reason:
Answer #1

Ans:

1)

2)n=27

df=27-1=26

critical chi square value=CHIINV(0.95,26)=15.379

Reject H0,if chi square<15.379

3)

Test statistic:

Chi square=(27-1)*12.3^2/14.1^2=19.785

4)p-value=CHIDIST(19.785,4)=0.1984

Do not Reject H0,as p-value>0.05

There is not sufficient evidence to support the claim that this class has a standard deviation less than 14.1